OpenStudy (anonymous):

In scientific notation, 3.0 x 10-5 + 4.0 x 10-4 =

OpenStudy (anonymous):

please help me

OpenStudy (anonymous):

I would Advise u to take 10^-5 common and then solve it

OpenStudy (anonymous):

If u have problem wid tht i will show u first step ... the rest shud be done by u if u take 10^-5 common then 10^-5( 3 + 4/10)

OpenStudy (anonymous):

3.0 time 10^-5 +4.0^-4 =?

OpenStudy (anonymous):

time 10^-4

OpenStudy (anonymous):

3 x 10^(-5) + 40 x 10^(-5) = 43 x 10^(-5)

OpenStudy (anonymous):

here 3.0 time + 10^-5 +4.0 time 10^-4 =?

OpenStudy (anonymous):

sorry I wrote wrong last time lol

OpenStudy (anonymous):

or 4.3 x 10^(-4)

OpenStudy (anonymous):

3.0time 10^-5 + 4.0 time 10^-4 =?

OpenStudy (anonymous):

yes how you solve it?

OpenStudy (anonymous):

\[3 \times 10^{-5} = 0.3 \times 10^{-4}\] \[0.3 \times 10^{-4} + 4 \times 10^{-4} = 10^{-4} \times (0.3 + 4) = 4.3 \times 10^{-4}\]

OpenStudy (anonymous):

ok I understand now

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

Welcome! :)